# TokMem: Tokenized Procedural Memory for Large Language Models

### v1

@@ 3.2 Atomic Memory Recall
Figure 2: Sample efficiency on a 10-task mixture from SNI. TokMem consistently outperforms fine-tuning in the low-data regime. TokMem can surpass RAG with only $10$ training samples, demonstrating strong few-shot learning capability.
@@
TokMem provides the most consistent and scalable performance across models and task scales. As shown in Table 1, non-parametric methods such as Base shows stable but fail to achieve competitive performance. RAG performs reasonably well on when memory is not heavy but quickly degrade as the number of task memory increases, indicating its sensitivity to retriever quality. Parametric methods such as fine-tuning achieve stronger initial accuracy but suffer from forgetting as tasks accumulate; replay memory alleviates this issue but still falls short of TokMem. By contrast, we see that TokMem maintains high accuracy with minimal performance drop when acquiring new task memories, achieving the best average results across all settings. The decoupled variant (TokMem+DC) provides modest gains for the smaller Qwen 0.5B model but provides no improvement for larger Llama models, and in some cases underperforms TokMem when scaling to many tasks.
@@
Table 2 further highlights TokMem’s robustness in memory routing. Its accuracy remains above $94\%$ even at 1,0001{,}000 tasks with the smallest 0.5B model, significantly outperforming the Sentence-BERT retriever used in RAG, whose accuracy drops below $80\%$ when have the stress to route for 1,0001{,}000 tasks. This high-fidelity memory routing enables TokMem to sustain strong performance without relying on external retrieval mechanisms or fine-tuning, demonstrating its advantage in continual and large-scale task acquisition.
Analysis of Training Efficiency.
@@
We compare the training sample efficiency of LoRA fine-tuning and TokMem on the first $10$ tasks from SNI, a mixture setup that removes the impact of forgetting. We set the adapter rank to $r=1$, which helps prevent overfitting and aligns its parameter scale with TokMem. The result in Figure 2 shows that TokMem consistently achieves higher performance than fine-tuning across all sample budgets, with the greatest advantage appearing in the low-data regime. The decoupled variant (TokMem+DC) offers small but consistent improvements, particularly when more samples are available. Overall, these results highlight TokMem’s ability to learn new procedures effectively with limited data, making it a both parameter and data efficient approach to memory acquisition.
3.3 Compositional Memory Recall
Dataset Details.
We construct a benchmark from the APIGen dataset (Liu et al., 2024b) by sampling the $50$ frequently used tools. Here, each tool invocation is treated as an atomic procedure, and solving a query requires composing multiple such procedures. We synthesize 5,0005{,}000 training queries and $500$ test queries, both capped at four calls. Details of the dataset can be found in Appendix D.2
We report performance using two F1 metrics: (i) Tool Prediction F1, which measures whether the correct tools are invoked; and (ii) Argument Generation F1, which evaluates the correctness of function call arguments. For robustness to semantic equivalence, both gold and predicted outputs are normalized into Abstract Syntax Trees before scoring (Patil et al., 2025).
Compositional Adaptation Fine-tuning.
We found that TokMem benefits from a brief adaptation phase that exposes the backbone with the compositional structures of memory tokens. Concretely, we fine-tine the backbone on a held-out auxiliary tool set using the same LoRA fine-tuning setup as the baseline. The adapted weights are then merged, after which the backbone remains frozen for memory acquisition and evaluation (see Appendix B.1 for details).
@@ 3.3 Compositional Memory Recall
Table 3: Compositional tool-use performance on APIGen. TokMem achieves strong tool selection and argument F1 across multiple calls, outperforming ICL and RAG with lower input-augmentation complexity, and surpassing fine-tuning with far fewer trainable parameters.
@@
Table 3 shows that TokMem achieves the strongest overall performance. Without adaptation for compositionality, it outperforms RAG while avoiding the added complexity from an external retrieval mechanism. Non-parametric baselines such as ICL and RAG perform poorly on both tool prediction and argument generation, particularly with the smaller Llama 1B model, likely due to its weak instruction-following ability.
Compared with parametric baselines, TokMem consistently matches or surpasses LoRA fine-tuning while requiring an order of magnitude fewer trainable parameters. Notably, TokMem exhibits stronger interpretability between tool selection and argument generation, with improvements in the former translating directly into the latter. By contrast, LoRA fine-tuning shows weaker alignment. For example, as seen with the 1B model, it often generates plausible arguments even when tool selection is incorrect, indicating that its argument generation is not properly grounded in the chosen tools.
Table 4: TokMem shows strong out-of-domain compositional generalization, significantly outperforming fine-tuning when tested on longer call chains than seen during training. Shaded cells indicate out-of-domain test scenarios.
Argument
Training Method 2 calls 3 calls 4 calls Avg.
[…]
3-call Fine-tuning 86.9 $5 \cellcolorbaseline8$0.3 84.2
[…]
4-call Fine-tuning 87.9 86.6 82.9 85.8
TokMem 85.9 86.7 88.3 86.3 (+0.5)
Compositional Generalization.
We observe that TokMem provides clear advantages in compositional generalization over fine-tuning. Table 4 reports Argument F1 when the Llama 3B model is evaluated on queries requiring more function calls than those observed during training.
Notably, when trained solely on single-call data, TokMem achieves much stronger performance than fine-tuning when evaluating on $2$ to $4$ calls test data. This demonstrates that memory tokens trained for atomic procedures can be effectively composed at test time, enabling strong zero-shot generalization to multi-step behavior.
As the training regime is expanded to include more calls (e.g., $3$ or $4$), the performance gap narrows, but TokMem remains competitive or slightly ahead across all configurations.
Figure 3: Forgetting analysis in continual adaptation. As new tools are introduced, fine-tuning with replay memory suffers sharp drops on earlier tasks, while TokMem maintains stable performance. Larger models show stronger retention due to greater capacity.
Figure 4: Effect of renormalization on TokMem. Without renormalization, new tokens dominate and older ones are forgotten, particularly in smaller models with limited embedding capacity. Renormalization effectively mitigates this by balancing norms across tokens.
@@
We compare TokMem with replay memory in a continual adaptation setting, where tools are introduced sequentially over five training rounds (e.g., tools 1–10 in the first round, 11–20 in the second, and so forth). As shown in Figure 3, we see that replay memory struggles to prevent catastrophic forgetting as new tools are introduced. By contrast, TokMem maintains higher performance across tool groups, with only mild declines that primarily reflect the growing number of tools. Larger models exhibit better retention for both approaches, likely due to their expanded parameter capacity, which reduces the risk of interference with previously learned tools.
We further investigate the effect of the renormalization step introduced in Section 2.3 on newly added memory tokens, whose norms may otherwise dominate older tokens in the softmax. As seen in Figure 4, TokMem without renormalization shows noticeable forgetting especially when the size of the model is small. However, larger models are more robust to forgetting even without renormalization, again due to their greater embedding capacity. Overall, renormalization improves TokMem’s resistance to forgetting by balancing routing between both new and old memory tokens. Additional analysis on the benefits of keeping the backbone frozen for continual memory acquisition is provided in Appendix B.2.
@@ 3.4 Analysis on Memory Placement
An important design choice in TokMem is the placement of memory tokens within the input sequence, which directly influences how the backbone model attends to and integrates procedural knowledge. While TokMem introduces a memory routing mechanism for generating tokens, its effectiveness also depends on this placement strategy. In the absence of routing, TokMem reduces to a prompt-tuning method (Li & Liang, 2021; Lester et al., 2021) with learnable embeddings, but it distinctively adopts an infix placement: query $\oplus$ MEM $\oplus$ response. Our experiments indicate that this infix design allows memory tokens to be activated after the context has been encoded, enabling context-aware conditioning and natural composition of multiple procedures.
However, it is unclear whether this memory placement is strictly better than the more common prefix formulation: MEM $\oplus$ query $\oplus$ response used in prior prompt-tuning work, where prefix tokens influence generation without having observed the query. To study the impact of placement, we compare prefix and infix placements under matched token budgets in the single-task setting.
Setup.
We compare TokMem with infix memory placement against prefix tuning by stress-testing the capacity of memory tokens using the recent Fanfics dataset collected after the pretraining of LLMs (Kuratov et al., 2025). We fix the sequence length to $128$ tokens and vary the batch size from $8$ to $32$, compressing batches of $1024$ to $4096$ response tokens into $1$ to $5$ memory tokens. For each sequence, we prepend a randomly generated query that serves only as a marker to distinguish the two placements, while the actual target to be learned remains the response.
We measure learning speed using Steps@90%Best, defined as the number of training steps (evaluated every 100 steps) required to reach 90% of the best perplexity. Results are averaged over five runs. Additional experiments evaluating generalization on a math reasoning dataset are provided in Appendix C.
Table 5: Comparison of TokMem vs. prefix tuning on memorizing text from the Fanfics dataset. TokMem converges faster and achieves lower perplexity than prefix tuning, particularly with few memory tokens.
@@ Results.
Table 5 shows that TokMem consistently achieves lower perplexity and often converges faster than prefix tuning. With a single token, TokMem reaches $90\%$ of the best perplexity roughly $30\%$ sooner than prefix tuning, indicating that conditioning memory after the query helps the model learn more efficiently. Interestingly, when more tokens are available (e.g., five tokens), the performance gap narrows. This suggests that prior work (Li & Liang, 2021), which typically uses dozens or even hundreds of tokens, may have underestimated the importance of memory placement in low-token regimes, where each token must compress more procedural information.
@@ 4 Related Work
Equipping LLMs with memory has been explored through multiple directions. Most existing approaches emphasize declarative memory, where the objective is to store and retrieve explicit information such as facts or conversation history (Packer et al., 2023; Chhikara et al., 2025; Zhong et al., 2024). In contrast, parameter-based approaches internalize task-specific behaviors within model parameters, resembling procedural memory. TokMem builds on this latter view while emphasizing modularity and compositionality.
@@
A common approach is to externalize memory as textual content retrieved at inference time. Retrieval-augmented generation (RAG)(Lewis et al., 2020) and its variants(Guu et al., 2020; Karpukhin et al., 2020; Borgeaud et al., 2022; Khandelwal et al., 2020) attach relevant textual chunks during inference, while RET-LLM (Modarressi et al., 2023) encodes knowledge as symbolic triplets. Building on these ideas, more recent systems such as MemGPT (Packer et al., 2023), Mem0 (Chhikara et al., 2025), and A-Mem (Xu et al., 2025) extend these ideas to conversational settings through hierarchical or summarization-based memory states. While effective for factual recall, these approaches are not optimized for procedural control and often incur significant inference-time overhead due to the re-read of textual memory.
@@
Another line of work encodes memory directly into model parameters. Fine-tuning and multitask instruction tuning (Wei et al., 2022a; Sanh et al., 2021), as well as parameter-efficient variants such as LoRA (Hu et al., 2022) allow models to acquire new procedures, but task knowledge is entangled. MemoryLLM (Zhong et al., 2024) introduces latent memory pools but remian entangled. Prompt-based methods such as prompt tuning (Lester et al., 2021; Wu et al., 2024; 2025) store knowledge implicitly as global embeddings without selective routing, and L2P (Wang et al., 2022b) introduces modular prompt pools but still relies on an external controller to determine which prompts are retrieved. ToolGen (Wang et al., 2025) compresses tools into virtual tokens but focuses on post-training the backbone through multi-stage fine-tuning. By contrast, TokMem keeps the backbone frozen, and introduces discrete, composable memory units that can be added or composed without retraining, supporting continual adaptation.
@@
A complementary direction explores how models compose skills from simpler building blocks. Chain-of-thought prompting (Wei et al., 2022b) and tool-augmented reasoning frameworks such as Toolformer (Schick et al., 2023) enable multi-step reasoning, but rely on textual instructions that must be re-interpreted at each step. Modular parameter methods (Rosenbaum et al., 2018; Pfeiffer et al., 2021) create specialized adapters that can be recombined, but composition requires parameter merging or heuristic routing. TokMem differs by representing procedures as discrete tokens that can be chained directly in context, enabling lightweight parameter-isolated composition.
5 Conclusion and Future Directions
@@
We introduced Tokenized Memory (TokMem), a parameter-efficient framework that encodes procedural memory as compact tokens. TokMem enables selective recall and compositional use of procedures without modifying backbone parameters, achieving strong performance across multitask and tool-augmented reasoning benchmarks.
Our experiments are conducted on the SNI and APIGen datasets, which allow controlled analysis of atomic and compositional recall. While these settings demonstrate the effectiveness of tokenized procedural memory without backbone training, they do not fully capture the diversity of real-world procedures. In particular, richer forms of composition, such as interleaving function calls with NLP tasks from SNI, as illustrated in Figure 1b, and multi-turn interactions remain unexplored, but could be supported with curated datasets. Overall, advancing TokMem toward practical deployment will require more realistic benchmarks or user-driven data collection pipelines that better reflect open-domain procedural knowledge.
Additional promising directions include incorporating reinforcement learning to improve generalization for complex compositional structures, and enabling personalization by allowing users to attach their own memory banks while keeping the backbone frozen. Together, these extensions pave the way for scalable, compact, and user-adaptive memory systems in large language models.
Appendix A Decoupled Embedding for TokMem

### v2

@@ 3.2 Atomic Memory Recall
As shown in Table 1, baseline methods such as Base are stable but fail to achieve competitive performance. RAG performs reasonably well when the memory load is small, but degrades quickly as the number of stored task memories increases, indicating its sensitivity to the retriever’s quality. Parametric methods such as fine-tuning achieve stronger initial accuracy but suffer from forgetting as tasks accumulate; replay memory alleviates this issue but still falls short of TokMem. By contrast, TokMem maintains high accuracy with minimal performance drop as it acquires new task memories, achieving the best average results across all settings.
We also include a variant where each memory token is decoupled into an address token and a steering token. This decoupling separates the roles of memory tokens and increases the capacity of TokMem. We refer to this variant as TokMem with decoupled embeddings (TokMem+DC); further details are provided in Appendix A.2.
The decoupled variant (TokMem+DC) yields modest gains for the smaller Qwen-0.5B model but no improvement for larger Llama models, and in some cases it underperforms TokMem when scaling to many tasks. Overall, although TokMem+DC is a tempting variant, it does not provide additional benefits. We therefore focus on the simple yet effective TokMem (without DC).
@@
Table 2 further highlights TokMem’s robustness in memory routing. Its accuracy remains above $94\%$ even at 1,0001{,}000 tasks with the smallest 0.5B model, significantly outperforming the Sentence-BERT retriever used in RAG. By contrast, the retriever’s accuracy drops below $80\%$ when routing over 1,0001{,}000 tasks. This high-fidelity routing enables TokMem to sustain strong performance without external retrieval or fine-tuning, demonstrating its advantage for continual and large-scale task acquisition.
Figure 2: Sample efficiency on a 10-task mixture from SNI. TokMem consistently outperforms fine-tuning in the low-data regime and can surpass RAG with only $10$ training samples, demonstrating strong few-shot capability.
Analysis of Training Sample Efficiency.
@@
We compare the training sample efficiency of LoRA fine-tuning and TokMem on $10$ SNI tasks. We vary the number of training samples per task from $10$ (few-shot) to $500$, and report performance across these sample budgets in Figure 2. We set the adapter rank to $r=1$, which helps reduce overfitting in the low-data regime and matches the parameter scale of TokMem. Results show that TokMem consistently outperforms fine-tuning across all budgets, with the largest gains at small sample sizes. The decoupled variant (TokMem+DC) yields modest gains in this controlled setting, mainly at higher sample budgets. Overall, these results highlight TokMem’s ability to acquire new procedures with limited data, making it both parameter- and data-efficient.
@@
Table 3: Compositional tool-use performance on APIGen. TokMem achieves strong tool selection and argument F1 across multiple tool calls, outperforming ICL and RAG while requiring less input augmentation, and surpassing fine-tuning with far fewer trainable parameters.
@@
3.3 Compositional Memory Recall
Dataset Details.
We construct a benchmark from the APIGen dataset (Liu et al., 2024b) by sampling $50$ frequently used tools. Each tool invocation is treated as an atomic procedure, and solving a query requires composing multiple such procedures. We synthesize 5,0005{,}000 training queries and $500$ test queries, each capped at four tool calls. Additional dataset details are provided in Appendix A.3.
We report performance using two F1 metrics: (1) Tool Prediction F1, which measures whether the model invokes the correct tools; and (2) Argument Generation F1, which evaluates the correctness of tool-call arguments. To account for semantic equivalence, both gold and predicted outputs are normalized into Abstract Syntax Trees before scoring (Patil et al., 2025).
Adaptation for Compositionality.
We find that TokMem benefits from a brief adaptation phase that exposes the backbone to the compositional structure of memory tokens. Concretely, we fine-tune the backbone on a held-out auxiliary tool set using the same LoRA setup as the baseline. We then merge the adapted weights, after which the backbone remains frozen for memory acquisition and evaluation (see Appendix A.4 for details). Note that the auxiliary tool set is different from the evaluation set, so this procedure does not violate our frozen-backbone setup.
This adaptation teaches the model how to interleave responses with memory tokens. We therefore treat it as part of TokMem for compositional tasks, and use TokMem with adaptation as the default configuration.
@@ 3.3 Compositional Memory Recall
Table 3 shows that TokMem achieves the strongest overall performance. Even without adaptation for compositionality, it outperforms RAG while avoiding the added complexity of an external retrieval mechanism. ICL and RAG perform poorly on both tool prediction and argument generation, especially with the smaller Llama 1B model, likely due to its weaker instruction-following ability.
TokMem consistently matches or exceeds LoRA fine-tuning’s performance while requiring an order of magnitude fewer trainable parameters. For example, when applied to Llama 8B, LoRA requires 3.413.41M trainable parameters, whereas TokMem needs only 0.20.2M and achieves higher performance.
Notably, TokMem exhibits stronger correlation between tool selection and argument generation: improvements in the former translate directly into gains in the latter. By contrast, LoRA fine-tuning shows weaker alignment. For example, on the 1B model, it often generates plausible arguments even when tool selection is incorrect, suggesting that argument generation is not properly grounded in the chosen tool.
TokMem also exhibits stronger compositional generalization than fine-tuning, successfully executing queries with more function calls than it has seen during training. Detailed results are provided in Appendix B.1.
Figure 3: Forgetting analysis under continual adaptation. As new tools are introduced, fine-tuning with replay memory suffers sharp performance drops on earlier tasks, whereas TokMem remains stable. Larger models exhibit stronger retention, likely due to their greater capacity.
Figure 4: Effect of renormalization on TokMem. Without renormalization, newly learned tokens dominate and older tokens are forgotten, particularly in smaller models with limited embedding capacity. Renormalization mitigates this by balancing token embedding norms.
@@
We compare TokMem with replay memory in a continual adaptation setting, where tools are introduced sequentially over five training rounds (e.g., tools 1–10 in the first round, 11–20 in the second, and so forth). As shown in Figure 3, replay memory struggles to prevent catastrophic forgetting as new tools are introduced. By contrast, TokMem maintains higher performance across tool groups, with only mild declines that mainly reflect the increasing number of tools. Larger models exhibit better retention for both approaches, likely due to their greater capacity, which reduces interference with previously learned tools.
We further investigate the effect of the renormalization step introduced in Section 2.4. Without renormalization, the norms of newly added memory tokens can dominate older tokens in the softmax (see Appendix B.2). As shown in Figure 4, TokMem without renormalization exhibits noticeable forgetting, especially for smaller models. Larger models are more robust even without renormalization, likely due to their greater embedding capacity. Overall, renormalization improves TokMem’s resistance to forgetting by balancing routing between new and old memory tokens. We provide further analysis in Appendix B.3 on how keeping the backbone frozen helps prevent forgetting during continual memory acquisition.
@@ 3.4 Analysis on Memory Placement
An important design choice in TokMem is where memory tokens are placed within the input sequence, as this directly affects how the backbone model attends to and integrates procedural knowledge. Although TokMem introduces a memory-routing mechanism for generating tokens, its effectiveness also depends on the placement strategy. Without routing, TokMem reduces to a prompt-tuning method (Li and Liang, 2021; Lester et al., 2021) with learnable embeddings; however, it adopts an infix placement: query $\oplus$ MEM $\oplus$ response. It remains unclear whether this placement is consistently better than the more common prefix formulation, MEM $\oplus$ query $\oplus$ response, used in prior prompt-tuning work, where prefix tokens influence generation before observing the query. To study the impact of placement, we compare prefix and infix placements under matched token budgets in the single-task setting.
Setups.
We compare TokMem with infix memory placement against prefix tuning by stress-testing the capacity of memory tokens (Sastre and Rosá, 2025) on the recent Fanfics dataset, which was collected after the pretraining of the LLMs (Kuratov et al., 2025). We fix the sequence length to $128$ tokens and vary the batch size from $8$ to $32$, compressing batches of $1024$ to $4096$ response tokens into $1$ to $5$ memory tokens. For each sequence, we prepend a randomly generated query that serves only as a marker to distinguish the two placements; the target to be learned remains the response.
We measure learning speed using Steps@90%Best, defined as the number of training steps (evaluated every 100 steps) required to reach 90% of the best perplexity. Results are averaged over five runs. Additional experiments on generalization to a math reasoning dataset are provided in Appendix B.4.
Table 4: Comparison of TokMem and prefix tuning for memorizing text from the Fanfics dataset, evaluated by perplexity. The suffix “-k” refers to the number of embeddings, and Steps@90%Best represents the training steps needed to reach 90% of the lowest achieved perplexity. TokMem converges faster and achieves lower perplexity than prefix tuning, especially with a small number of memory tokens.
@@ Results.
Table 4 shows that TokMem consistently achieves lower perplexity and often converges faster than prefix tuning. With a single token, TokMem reaches $90\%$ of its best perplexity roughly $30\%$ sooner than prefix tuning, suggesting that conditioning on memory after the query helps the model learn more efficiently. Interestingly, when more tokens are available (e.g., five tokens), the performance gap narrows. This suggests that prior work (Li and Liang, 2021), which typically uses dozens or even hundreds of tokens, may have underestimated the importance of memory placement in low-token regimes, where each token must compress more procedural information.
@@ 4 Related Work
Equipping LLMs with memory has been explored along multiple directions. Most existing approaches emphasize declarative memory, where the objective is to store and retrieve explicit information such as facts and conversation history (Packer et al., 2024; Chhikara et al., 2025; Wang et al., 2024). By contrast, parameter-based approaches internalize task-specific behaviors within model parameters, resembling procedural memory. TokMem builds on this latter view while emphasizing modularity and compositionality.
@@
A common approach is to externalize memory as textual content retrieved at inference time. Retrieval-augmented generation (RAG; Lewis et al., 2020) and its variants (Guu et al., 2020; Karpukhin et al., 2020; Borgeaud et al., 2022; Khandelwal et al., 2020) attach relevant textual chunks during inference, while RET-LLM (Modarressi et al., 2024) encodes knowledge as symbolic triplets. Building on these ideas, more recent systems extend these ideas to conversational settings through hierarchical or summarization-based memory states; examples include MemGPT (Packer et al., 2024), Mem0 (Chhikara et al., 2025), and A-Mem (Xu et al., 2025). While effective for factual recall, these approaches are not optimized for procedural control and often incur significant inference-time overhead due to repeatedly re-reading textual memory.
@@
Another line of work encodes memory directly into model parameters. Fine-tuning and multitask instruction tuning (Wei et al., 2022a; Sanh et al., 2022) allow models to acquire new procedures, but task knowledge often becomes entangled within the shared weights. Parameter-efficient fine-tuning methods such as LoRA (Hu et al., 2022) and Flora (Hao et al., 2024) disentangle new knowledge from the pre-trained base model, but still suffer from interference when learning multiple tasks. Mixture-of-LoRAs (Feng et al., 2024) address this entanglement issue by dynamically routing inputs to task-specific experts; but the mixtures are typically invoked independently and are not designed for memory composition. MemoryLLM (Wang et al., 2024) introduces latent memory pools but remains entangled. Prompt-based methods such as prompt tuning (Lester et al., 2021; Wu et al., 2024; 2026) store knowledge implicitly as global embeddings without selective routing, and L2P (Wang et al., 2022b) introduces modular prompt pools but still relies on an external controller to determine which prompts are retrieved. Prompt compression methods (Mu et al., 2023; Chevalier et al., 2023) compress prompts into context-agnostic representations, which may distort prompt information. ToolGen (Wang et al., 2025) compresses tools into virtual tokens but focuses on post-training the backbone through multi-stage fine-tuning. By contrast, our TokMem keeps the backbone frozen and introduces dedicated memory embeddings that can be added or composed without retraining, supporting continual adaptation.
@@
A complementary direction explores how models compose skills from simpler building blocks. Chain-of-thought prompting (Wei et al., 2022b) and tool-augmented reasoning frameworks such as Toolformer (Schick et al., 2023) enable multi-step reasoning, but they rely on textual instructions that must be re-interpreted at each step. Modular parameter methods (Rosenbaum et al., 2018; Pfeiffer et al., 2021) create specialized adapters that can be recombined, but composition typically requires parameter merging or heuristic routing. TokMem differs by representing procedures as discrete tokens that can be chained directly in generation, enabling lightweight, parameter-isolated composition.
5 Conclusion
@@
We introduced TokMem, a parameter-efficient framework that encodes procedural memory as compact tokens. TokMem enables selective recall and compositional use of procedures without modifying backbone parameters, achieving strong performance across multitask and tool-augmented reasoning benchmarks.
Future directions are discussed in Appendix C, including the automation of query–procedural decomposition for synthesizing training trajectories, reinforcement learning for stronger compositional generalization and personalization through user-specific memory banks. These extensions pave the way for scalable, compact, and user-adaptive memory systems in large language models.
Acknowledgments
We thank the reviewers and chairs for their efforts. The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Amii Fellow Program, the Canada CIFAR AI Chair Program, a donation from DeepMind, and the Digital Research Alliance of Canada (alliancecan.ca).
Appendix A Experimental Details
A.1 Details of Super-Natural Instructions
We sample 1,0001{,}000 English tasks from the SNI dataset, where each task is labeled with a task ID and a short descriptive name. The full list of sampled tasks is provided in Figure 8. Tasks are introduced to the model sequentially in ascending order of their IDs, for instance, the model first sees task 1, then task 2, and so on.
After training on the first $k$ tasks, we save a checkpoint and evaluate performance on the test set of each of the $k$ tasks encountered so far. This simulates a continual learning setup where the model is expected to acquire new procedures while retaining previously learned ones. Once the model has been trained on all 1,0001{,}000 tasks, it is expected be able to perform all of them without forgetting earlier tasks.
A.2 Details of Decoupled Embedding for TokMem